What value would be needed to complete the following probability distribution?
0 1/3 = 8/24
1 1/8 = 3/24
2 1/8 = 3/24
3
4 1/6 = 4/24
Missing 6/24 = 1/4
I'm a little confused
Could you show me how they got the missing number and the 8/24, 3/24, 3/24, 4/24..?

Answer by stanbon(75887)

What value would be needed to complete the following probability distribution?
0 1/3 = 8/24
1 1/8 = 3/24
2 1/8 = 3/24
3
4 1/6 = 4/24
Missing 6/24 = 1/4
I'm a little confused
Could you show me how they got the missing number and the 8/24, 3/24, 3/24, 4/24..?
-------
The probabilities must add up to 1.
(8/24+3/24+3/24+4/24) + x = 1
18/24 + x = 1
x = (24/24)-(18/24) = 6/24 = 1/4
---------------
Cheers,
Stan H.
-----------